Simon is 26 years old and weighs 154 pounds.  A year ago, he decided to stop eating meat because he thought it would make his diet healthier.  Now that he is studying nutrition, he has become concerned that his vegetarian diet may not be as healthy as he though.  Simon records his food intake for one day and uses an online nutrient analysis program to assess his nutrient intake.  His analysis reveals that his diet provides 2900 calories, 78 g of protein, and 43 g of saturated fat.  

Does Simon’s diet provide enough protein to meet his RDA of 0.8 g of protein per kilogram of body?

How does the percentage of calories from saturated fat in Simon’s diet compare with recommendations?
Answer #1

1. Weight of Simon = 154lbs = 154×0.45

= 69.3 kg

Age = 26years

According to his analysis his diet provides :-

Calories - 2900calories

Protein - 78gm

Saturated fat = 43gm

According to RDA ,his protein intake should be 0.8× 69.3

= 55.44gm ,so his diet is providing sufficient protein ,even more than his requirements

2. 1gm of fat provides 9kcal

According to his analysis ,his diet is providing 43gm of saturated fat

43gm of fat will provide 43×9 = 387 calories

His whole diet provides 2900 calories

% of calories provided by fat = 387 / 2900 × 100

= 13.3 % of calories is provided by saturated fat

According to the latest Dietary guidelines for Americans recommends keeping saturated fat to less than 10 percent of calories a day . So , Simon should decrease his fat intake .